### v1

# Dialectal and Low Resource Machine Translation for Aromanian

Abstract
We present a neural machine translation system that can translate between Romanian, English, and Aromanian (an endangered Eastern Romance language); the first of its kind. BLEU scores range from 17 to 32 depending on the direction and genre of the text. Alongside, we release the biggest known Aromanian-Romanian bilingual corpus, consisting of 79k cleaned sentence pairs. Additional tools such as an agnostic sentence embedder (used for both text mining and automatic evaluation) and a diacritics converter are also presented. We publicly release our findings and models.111https://huggingface.co/aronlp Finally, we describe the deployment of our quantized model at https://arotranslate.com.
[…]
Dialectal and Low Resource Machine Translation for Aromanian
Alexandru-Iulius Jerpelea Alina-Ștefania Rădoi Sergiu Nisioi Tudor Vianu National College of Computer Science, Bucharest West University of Timișoara Human Language Technologies Research Center, University of Bucharest alex.jerpelea@gmail.com, alina.radoi@e-uvt.ro, sergiu.nisioi@unibuc.ro
1 Introduction
Training good machine translation systems in the low-resource setting is a task far from being solved Wang et al. (2021); Haddow et al. (2022). Current advances in large language models and evaluation methodologies are centered on English or are massively multilingual, but do not engage with the particularities of low-resource languages. Of the 7000+ languages spoken in the world, only a small part is covered by current machine translation (MT) systems Costa-jussà et al. (2022).
Aromanian (ISO 639-3 - rup) is an endangered Campbell et al. (2022) Eastern Romance language that currently lacks large-scale corpora and electronic resources that can potentially contribute to the preservation of its cultural heritage. In this study, we are interested in building a dataset suitable for training machine translation systems in the low-resource setting between two closely related languages from the same dialectal continuum: Romanian (ro) and Aromanian (ISO 639-3 - rup). Given the close similarity Caragiu-Marioțeanu (1975) between the two languages, we hope to expand the current pretrained LLM and machine translation models available for standard Romanian to its Eastern Romance sibling, in order to contribute to the preservation of this endangered222https://endangeredlanguages.com/lang/963, Accessed: 2024-09-16 language.
We believe the difficulties for such a task stem from the fact that: (1) there is little availability of monolingual or bilingual texts, (2) the writing system, vocabulary, and grammar have not been standardized in a widely accepted institutional manner, and several types of spelling are currently being used Caragiu-Marioțeanu (1997); Cunia (1997); Nevaci (2008); Ballamaci (2010), and (3) Aromanian has several varieties / dialects that have been influenced by the contact with Greek, Romanian, Turkish, Albanian, and South Slavic languages Caragiu-Marioțeanu (1975). Furthermore, the language has historically been transmitted orally within families or small communities Gica et al. (2009); Maiden (2016), with public usage limited to a few localities in Albania and North Macedonia.
Aiming to fill this gap, our contribution can be summarized as follows:
the creation of a 79k multigenre parallel Aromanian-Romanian sentence-aligned parallel corpus, augmented with with machine translated English sentences
a LaBSE sentence encoder Feng et al. (2020), fine-tuned for Aromanian support
a comparison of different NLLB-200 Costa-jussà et al. (2022) models fine-tuned to translate in any direction between Aromanian, Romanian, and English.
a range of instruction-tuned large language models (LLMs), fine-tuned for translation
a diacritics converter between the multiple orthographic variations of Aromanian
2 Related Work
Low resource MT has gained considerable traction in the past few years. Initiatives range from commercial ambitions to build ”machine translation systems for the next thousand languages” Bapna et al. (2022), the NLLB family of models that can translate between more than 200 languages Costa-jussà et al. (2022), to approaches that leverage generative language models Alves et al. (2024); Xu et al. (2023). Options for closely related languages employing transliteration or transformer models have been attempted for Belorussian to Russian and Arabic dialects Karakanta et al. (2018); Kumar et al. (2021). To our knowledge, no successful attempt has been made to integrate Aromanian in a translation system.
In terms of electronic resources for Aromanian, they consist mainly of multilingual word-aligned lists (Nisioi, 2014; Cristea and Dinu, 2020; Beniamine et al., 2020; Fourrier and Sagot, 2022; Mititelu et al., 2021) to study language evolution, language contact, historical linguistics, and the relationship with other Romance languages.
The recent work of Petrariu and Nisioi (2024) presents a multilingual parallel corpus of approximately 3k sentences covering Aromanian. Although this is the largest published resource so far, the corpus only covers one genre, namely fairy tales, and the total number of sentence pairs is relatively small, rendering it insufficient to train qualitative MT models, as concluded by Petrariu and Nisioi (2024) when obtaining a maximum of 8 BLEU points.
Our work is the first to build an end-to-end machine translation system for Aromanian language together with the release of a large multi-genre corpus.
3 Collecting a Multigenre Dataset
In this section, we describe the extraction and alignment method for each parallel text used to train and test our MT system. These texts pertain to different genres such as news articles, literature, dictionaries, religious texts, music lyrics, and essays. They focus on the Romanian-Aromanian language pair.
The Bible has been translated into Aromanian by Dina Cuvata and was published in 2004. The only available version that the we could find is an online scan of the printed edition.333https://dinitrandu.com/wp-content/uploads/2022/04/Bibliea-limba-armaneasca.pdf, Accessed: 2024-09-16 We manually check each page of the scan and perform OCR with Tesseract V5 Smith (2007) on the pages that are intelligible. Aromanian is not supported by the engine, so we set the language parameter to Romanian. We then carefully check each verse and pair it with its match from the Romanian translated Bible. Each verse is divided into sentences in both languages. If the number of sentences per verse matches, we pair the sentences and add them to our corpus. We obtain around k sentence pairs from this source.
Lyrics Translate444https://lyricstranslate.com, Accesed: 2024-09-16 is a platform where songs are translated into multiple languages. We scrape the 500+ songs in both Aromanian and Romanian versions, and pair them verse by verse, obtaining k pairs.
The Multilingual Parallel Corpus of Aromanian (MPC-rup) consists of approximately 3000 pairs of sentences from Aromanian fairy tales and short prose Petrariu and Nisioi (2024).
A list of parallel short phrases and a set of anonymously authored long sentences and phrases dictionary, consisting of k pairs, which we have received as a donation.
The Avdhela project,555https://www.proiectavdela.ro Accessed: 2024-09-16 is a digital Aromanian library consisting of a collection of parallel poetry and prose. We extract 4000 pairs.
Radio Romania International is a Romanian public news radio station that includes editorials in many languages, among them Aromanian. The majority of radio transmission are published in text form on the official website666https://www.rri.ro/ro_ar, Accessed: 2024-09-16 and Aromanian articles are usually translations of matching Romanian articles.777A fact confirmed by several authors working for the radio station. However, there are no linkbacks to these pages, which makes it difficult to trace back the corresponding articles. Thus, we scrape all articles in both languages and match them by the images they contain, as we believe this to be an efficient proxy. In cases where multiple matches are possible, we deploy our sentence alignment tool (described in section 4) between the set of titles in Aromanian and the set of titles in Romanian, pairing the titles with the greatest semantic similarity, and retaining the possibility of unmatched titles. This method yields k sentence pairs.
The adventures of Tom Sawyer Twain (2007) was translated into Aromanian from Romanian and donated to us (as it is not yet published). We align the sentences with the same alignment tool described in section 4 and yield a total of k sentences.
A century of Aromanian poetry Cândroveanu and Iorgoveanu (1985) is a collection of Aromanian poetry translated into Romanian. We use noisy scanned version and the Open CV Bradski (2000) computer vision tool to identify text bounding boxes (Figure 1), and, considering their location within the image, establish if the text chunk is in Aromanian or Romanian. Next, for each page, we align the resulting sentences using the alignment tool described in section 4, which adds k pairs to the corpus.
The Divine Comedy Dante (2017) was translated by Dina Cuvata in Aromanian888https://dinitrandu.com/wp-content/uploads/2018/08/Dante-Dina-Cuvata.pdf, Accessed: 2024-09-16, and by aligning it with the Romanian version we obtain k pairs.
An anonymously donated collection of modern Aromanian poetry, which cannot be found online and is already aligned, contains k pairs.
Tao Te Ching has been partially translated999https://dinitrandu.com/wp-content/uploads/2022/06/Cartea-a-Calillei.pdf, Accessed: 2024-09-16. By aligning it with its Romanian counterpart, we obtain 260 sentences.
The translated Aromanian version of ”The Little Prince” is aligned similarly to the Romanian version with sentence pairs.
Writings is a collection of news articles consisting of news articles, personal stories, and essays, which are not available online. These were donated to us by Kira Mantsu, a prominent Aromanian writer. The documents are originally in Aromanian and translated into Romanian by the author herself; the alignment process results in a total of k new pairs.
3.1 Dataset Split
Considering all the sources mentioned above, the dataset we use to build our models consists of k sentence pairs. A table with sample sentences from each source can be viewed in Appendix A.
In preparation for the model training process, we split our dataset into train, dev and test splits. Sentences from ”The Little Prince” and Writings are only included in the dev and test sets, but are not used for training. The reason behind this choice is that we want to evaluate the model’s performance on novel data, and we believe that the literary genre is a difficult evaluation task.
For our train set, we extract in a stratified manner random pairs from each of the above sources. The remaining pairs from those sources make up our dev set. We also add the sentences from ”The Little Prince” to our dev set to observe model evaluation scores during training. The test set comprises only Writings documents to enforce evaluation on out-of-domain data; thus, the texts do not have a common authorial origin with the data on which our models are trained. The total number of sentence pairs per source per split can be visualized in Table 1.
| Split | Bible | Lyrics | Small corpus | Phrases | Avdhela | Radio | Poetry | Tom Sawyer |
| train | ||||||||
| dev | ||||||||
| test |
| Split | Divine Comedy | Modern Poetry | Tao Te Ching | The Little Prince | Writings | Total |
| train | ||||||
| dev | ||||||
| test |
3.2 Synthetic Data
Because our aim is to also perform translations between Aromanian and English, we add synthetic English counterparts to all of the sentence pairs by translating Romanian sentences with the Google Translate API.
To evaluate the quality of automatic English translations, we employ the Cometkiwi-da model Rei et al. (2022) for reference-free evaluations, therefore English translations are evaluated only with regard to the Romanian original sentences. Throughout the corpus, we obtain a system-score of (scores range from to ). The resulted score is an indicator that Romanian-to-English machine translations are of mediocre quality and that automatic translations to and from English are error-prone, but still meaningful.
3.3 Dataset Size
It is important to note that there are numerous orthographic standards for Aromanian. The main spelling types in our corpus are DIARO (named after the Aromanian-Romanian Dictionary by Caragiu-Marioțeanu (1997), and Cunia, named after Tiberius Cunia, the author of the Aromanian Dictionary, Cunia (2010). We also sparsely found Greek orthography, which we standardized using the Latin script.
The Cunia and DIARO spellings are easily convertible to each other using simple regular expressions, with the exception of the closed central and mid-central vowels. Both of these sounds are represented by the letter ¡ã¿ in the Cunia standard. However, the DIARO standard follows the Romanian standardization Petrariu and Nisioi (2024) using two different symbols ¡î¿ and ¡â¿ for closed central and ¡ă¿ for mid central.
We chose to convert our entire corpus to the Cunia writing system, the main reason for this choice being the slightly lower fertility rate of the BPE tokenizers versus DIARO orthography that reaches tokens per word. We observed lower fertility scores across various models including both NLLB-200 and pre-trained large language models. Since the pre-trained models are multilingual, it is likely that the tokenizers have seen byte-pairs similar to Cunia orthography in words from Albanian, South Slavic, and Greek Latin transliterated versions.
Combating the potential loss of details due to the merging of the mid-central and closed central vowels, we train an n-gram-based model to convert from Cunia to DIARO.
Table 2 presents several statistics regarding our Cunia-converted datasets. Original Aromanian texts have the greatest lexical richness and a tendency to create shorter phrases. However, automatically aligned human translations into Romanian and the machine translated texts in English have both lower lexical diversity (as estimated by type-token ratio) and larger sentence lengths. In Appendix B, we also present these statistics for each text genre separately.
| language | words | unique words | type-token ratio | words/sentence |
| Aromanian | 1,189,000 | 175,000 | 0.15 | 15.31 |
| Romanian | 1,279,000 | 114,000 | 0.09 | 16.45 |
| English | 1,371,000 | 68,000 | 0.05 | 17.64 |
4 Aromanian Sentence Embeddings
Throughout the data collection process, we used a sentence embedder compatible with Aromanian to align documents. Thus, we fine-tune a language-agnostic Bert Sentence Encoder (LaBSE) Feng et al. (2020) model with an extended vocabulary, using a similar approach to Dale (2022). The purpose is to embed Aromanian sentences in the same latent space as the sentences from Romanian and English.
First, we extend the tokenizer’s vocabulary on a monolingual Aromanian corpus (i.e., we utilize Aromanian sentences from an intermediate version of the corpus). We perform this extension with the BPE algorithm Sennrich (2015), and we obtain k new tokens. This lowers the tokens per word ratio of LaBSE tokenizers from to .
Secondly, we fine-tune the model on both Aromanian-Romanian and Aromanian-English parallel pairs, updating only the embedding matrix, with the contrastive loss, in a similar manner to Feng et al. (2020).
At each training step, we randomly choose a batch of pairs in a random pair of languages out of the two established language pairs. We calculate the dot product of every possible sentence pair’s embeddings and only reward the pairs that are supposed to match. However, we penalize the matching pairs by substracting a small margin , to avoid over-fitting. We stop after 150k steps, as the graphs of both loss and accuracy (of matching translated sentences) seem to flatten after this point (Figure 2). Thus, in a batch of random sentence pairs, the model pairs each Aromanian sentence with its corresponding translation with an accuracy of more than . The training hyper-parameters are available in Appendix C.
The model can be used in multiple ways: (1) to calculate the BERTScore Zhang et al. (2019) and validate our MT models (see subsection 6.1); and (2) to align sentences in parallel documents. Similarly to Dale (2022), we use dynamic programming to select a sequence of sentence pairs that have the highest possible sum of sentence similarity scores. This method has been used to mine parallel sentences in section 3 from various genres.
5 Machine Translation Models
5.1 GPT Baseline
Inspired by recent MT results Hendy et al. (2023); Kocmi et al. (2024) where closed-source systems achieved state-of-the art performance on machine translation, we employ the GPT-4o model, in hopes of leveraging its extensive multilingual pre-training Achiam et al. (2023). Additionally, we suspect that the model had been exposed to Aromanian in its training stage. For example, the Aromanian Wikipedia101010https://roa-rup.wikipedia.org, Accessed: 2024-09-16 although of low quality, is part of the Common Crawl111111https://commoncrawl.org, Accessed: 2024-09-16. We use OpenAI’s gpt-4o model in zero-shot mode to translate all pairs in our test split, which we later evaluate in section 6.
| System prompt | You are a helpful assistant capable of accurately translating between Aromanian, Romanian, and English. |
| Instruction | Translate this sentence from Aromanian to English |
| Input | Shi-lji intrã Enoh tu vrearea-al Dumnidzã sh-ma multu nu s-aflã cã-l mutã Dumnidza. |
| Output | And God pleased Enoch, and then he was no more, because God moved him. |
| model | ron-¿rup | rup-¿ron | eng-¿rup | rup-¿eng | eng-¿ron | ron-¿eng |
| GPT-4o | 19.0 | 37.2 | 17.6 | 36.7 | 65.7 | 79.2 |
| Qwen2 7B Instruct | 42.1 | 48.9 | 35.1 | 50.9 | 56.6 | 75.7 |
| RoLLaMA | 42.1 | 48.9 | 35.9 | 51.3 | 58.4 | 76.9 |
| TowerInstruct v0.2 | 41.9 | 49.1 | 27.9 | 43.2 | 48.6 | 74.8 |
| LLaMA 3.1 8B Instruct | 42.4 | 49.6 | 35.2 | 51.7 | 58.9 | 76.9 |
| NLLB 1.3B | 43.7 | 51.5 | 39.5 | 53.9 | 61.5 | 75.9 |
| NLLB 600M | 45.0 | 51.0 | 40.4 | 53.0 | 60.1 | 75.1 |
5.2 NLLB Machine Translation
Seeking to benefit from transfer learning, we aim to fine-tune NLLB-200 Costa-jussà et al. (2022), an encoder-decoder transformer architecture model that can translate between any of its 202 languages. Our final goal is a model that has the ability to translate between Aromanian, Romanian, and English. The last two are already supported by the NLLB.
The NLLB tokenizer uses language tags, i.e. special tokens are added to the source and target texts, used by the model in the pre-training phase for identifying the source and target languages. The special tokens for Romanian and English are <ron_Latn> and <eng_Latn> respectively. We expand the tokenizer vocabulary with the <rup_Latn> token for Aromanian. Its meaning is supposedly constructed during the fine-tuning step. To achieve this, we initialize the embedding value of the newly added token as the mean of the <ron_Latn> and <ell_Grek> (Greek) embedding vectors. This is justified by the close relationship of Aromanian with standard Romanian and the large proportion of Greek influence and loanwords Dragomirescu (2020).
We experiment with the 1.3B and 600M distilled models of the original 3.3 billion parameter model. We train for k steps, where each sentence pair in the batch is in a random direction between Aromanian, Romanian, and English (i.e. six possible directions), aiming to better stabilize the training process this way. Also note that NLLB can already translate between Romanian and English. The hyper-parameters used for training the two models are available in Appendix D.
For later testing (section 6), we select the NLLB-600M checkpoint at k training steps, and the NLLB-1.3B checkpoint at k training, as these produce the highest BLEU scores on the dev split.
5.3 LLMs for Machine Translation
LLMs have proven useful in numerous NLP tasks, including machine translation. Ranging from zero-shot or few-shot translation Kocmi et al. (2024), to fine-tuning LLMs for MT Alves et al. (2024).
We choose to perform full fine-tuning on four instruction-tuned LLMs (i.e., base LLMs fine-tuned for instruction following). The models are as follows:
LLaMA 3.1 8B Instruct Dubey et al. (2024) - the most recent multilingual LLaMA model covering a diverse set of instructions; Romanian appears to be incidentally covered in the 15% of multilingual tokens given during pretraining
Qwen2 7B Instruct Yang et al. (2024) - a multilingual model from Alibaba Cloud covering as many as 30 languages, including 4 well resourced Western and Italo-Romance languages: French, Spanish, Portuguese, Italian
RoLLaMA 3 8B Instruct Masala et al. (2024) - a LLaMA3-based model fine-tuned specifically to respond to tasks and instructions in Romanian; the exposure of this model to Romanian should allow for an easier transfer to Aromanian, its Eastern Romance sibling
TowerInstruct 7B v0.2 Alves et al. (2024) - a LLaMA2-based model trained specifically to solve translation tasks (document-level translation, terminology-level translation, etc.) on 10 languages, including 4 well resourced Western and Italo-Romance languages: French, Spanish, Portuguese, Italian
Except for the RoLLaMA model, which has been specifically fine-tuned for Romanian language tasks, all the other models support Romanian only incidentally.
As with NLLB, we have the same objective of translating between Aromanian, Romanian, and English, so we train the LLMs with samples from all six possible directions. In the case of the four models, fine-tuning for more than epoch led to an increased loss on the dev split. Thus, we keep only the checkpoints fine-tuned for epoch with sample packing Krell et al. (2021). An example of a prompt for the LLaMA 3.1 model can be found in Table 3. More details about each model’s training hyper-parameters and prompt format can be found in Appendix E.
6 Evaluation
6.1 Automatic evaluation
| Participant | Direction | Fluency | Style | Meaning | |||
| HT | MT | HT | MT | HT | MT | ||
| Subject 1 | ron->rup | 9.66 | 9.73 | 9.60 | 9.60 | 9.73 | 9.86 |
| rup->ron | 9.46 | 9.66 | 9.46 | 9.53 | 9.46 | 9.33 | |
| Subject 2 | ron->rup | 9.66 | 9.66 | 9.33 | 9.60 | 9.66 | 9.73 |
| rup->ron | 8.13 | 9.0 | 7.73 | 8.73 | 8.2 | 8.93 | |
We evaluated all our models in all possible directions between Aromanian, Romanian, and English. We note here that the English references are machine translated and the results involving English should be taken with reservation. However, we provide the evaluation scores here for completeness.
The automatic metrics that we employ are BLEU Papineni et al. (2002), ChrF++ Popović (2015), and BERTScore Zhang et al. (2019). We report the ChrF++ scores for all the models in Table 4. All other results are available in Appendix F.
Overall, for English-Romanian pairs, GPT-4o appears to perform better than our models trained for the same language pairs. This could be an indicator that GPT-4o generates similar results to Google Translate, however, we cannot derive any conclusions at this point with respect to quality.
More importantly, the scores for the Aromanian-Romanian language pairs obtained by our models are considerably higher than those obtained by the GPT-4o model. Among our entire model set, the best performing are the NLLB family, obtaining consistently higher scores than their fine-tuned instruction LLM counterparts, regardless of the automatic metric (Appendix F). When translating into Aromanian (-¿rup), the differences are considerable - in every single case NLLB-type models perform significantly better than LLM-based translation models. However, when translating into a well-resourced language such as English or Romanian, the differences between NLLB models and LLM are less significant, only a few points apart.
All types of models struggle more to generate good quality Aromanian, a fact visible also in the overall lower BLEU scores (max 17 ron-¿rup, 13.4 eng-¿rup). A manual inspection of the translations reveals that the models tend to translate sentences word-by-word which misses word overlaps in BLEU-based metrics for out-of-domain evaluation on mixed-genre literary, news, and essay. More details on manual evaluation are presented in the next section.
6.2 Human Evaluation
Two human evaluators were involved in this study. Both were native Aromanian speakers residing in Romania, so they can speak both Romanian and Aromanian. We do not evaluate for English here. The sentences given for evaluation contain both human and machine translations in equal proportion and so each annotator evaluates a total of samples. For machine translation, we use our NLLB 600M fine-tuned model.
Annotators are instructed to assign direct assessment scores Graham et al. (2013) from 1 to 10 to evaluate the quality of each translation (1 being the lowest score, 10 being the highest). This scale is culturally motivated by the fact that it is familiar to the study participants because it is used in Romanian schools at all levels of post-primary school.
They assigned a score taking into account three categories: fluency, style, and meaning (logical sense). Fluency refers to how grammatically correct the sentence is. Style denotes how likely it is for the speaker to build the sentence they are evaluating. Meaning alludes to the logical sense of the sentence.
The evaluation was carried out bilingually for all language pairs. The results can be seen in Table 5. The outcome is dubious, as it may appear that machine translation is rated better than human translation. Manually analyzing the samples, we observe that oftentimes, the human translations are more creative, whilst the model tends to keep the structure of the original sentence mot-a-mot. In addition, the small number of human evaluators together with the small sample size renders the human evaluation inconclusive.
7 Online translation system
Seeking to make our machine translation model accessible to the public and raise awareness of the endangered status of the Aromanian language, we deploy a translation system online. Our GUI is lightweight, easy to use, and multilingual, with implemented accessibility features for copying text content and switching between source and target language (Figure 3).
| ron->rup | rup->ron | |||
| Model | BLEU | ChrF++ | BLEU | ChrF++ |
| Unquantized | 17.3 | 45.03 | 30.95 | 51.01 |
| Quantized | 17.01 | 44.84 | 30.21 | 50.16 |
7.1 Quantization
We choose to deploy on bare-metal CPU the 600M distilled NLLB version of our translator, as it scores best on a range of categories and has the smallest number of parameters. We quantize it using the ctranslate2 Klein et al. (2020) engine to bit-integer weights. Additionally, the engine also applies other optimizations such as layer fusion, batch reordering, etc.
Regarding the performance loss that is expected with quantization, we report minimal differences in the automatic metrics we deploy. These effects can be seen in Table 6 where we analyze the Aromanian-Romanian direction of our test split. Differences of less than BLEU point are measured.
Because all our models were fine-tuned for sentence-level translations, we split each user’s input into sentences and process it in parallel. Our system can translate approximately tokens per second.
7.2 Orthography Converter
[…]
8 Conclusions
Building a dialectal machine translation system for a low-resource language comes with several challenges - from the lack of standardization and small online presence of the language, to the diverse varieties specific to different countries. Given Aromanian’s close resemblance to Romanian, we show that it is possible to use pre-existing knowledge of Romanian both from LLM-like transformers and sequence-to-sequence models to build a MT system for Aromanian, Romanian, and English. Based on automatic evaluation, it appears that smaller Romanian-pre-trained seq-to-seq models are better translators than LLM-based English-centric or multilingual fine-tuned models (that have not been pre-trained on Romanian). Regarding human evaluation, due to the relatively small size of the community and lack of access to a large pool of native speakers, we had difficulties carrying out an extensive human evaluation. Future studies are mandatory to quantify the limitations of our MT system across different dialectal regions. With our multigenre dataset, we hope to bring Aromanian into the spotlight of MT and corpus-based linguistic research and contribute to preservation efforts for this endangered language.
9 Limitations
Aromanian is a very low resource language and is not standardized, having several varieties. Due to this, translation quality is experimental and oftentimes sub-par what a user might expect from a translation system for medium to high resource languages. This model is biased to produce the Aromanian variety spoken mostly in Romania. In addition, translation errors may occur where certain information is missed or altered. The words generated by the model might not be representative for speakers across all the regions and varieties where Aromanian is spoken.
Last but not least, this work mentions already a limitation in terms of lack of availablity to carry an in-depth speaker-centric human evaluation of the machine translation system. This is an issue we are looking into for an extended version of this work.
10 Ethical Considerations
We were able to collect and compile parallel translated texts from multiple sources. Note that for each text source, we have received explicit permission to use the text, the work is licensed in such a way that we are legally allowed to utilize and release it, or our scope falls under the fair use case.
11 Acknowledgements
This work would have not been possible without support from the Aromanian community, who helped us with the data collection processs, system testing, and gave us invaluable advice. We also acknowledge the support of the Human Language Technologies Research Center (University of Bucharest), for model training compute power, and hosting our online translation system.
Appendix A Aromanian-Romanian sample pairs per source
| Source | Aromanian | Romanian |
| Bible |
Shi vasiljelu Ptolometi muri dupã trei dzali, shi-ashchirladzlji cari eara pit tsitãts lj-vatamara oaminjlji di-a loclui.
Ti-agiutarea-a tsilor far di mindueari-arauda unã giudicata cu itsralji, a tinirlui om ti-unã cãnushteari shi bunã mindueari. |
Și regele Ptolomeu a murit a treia zi, și ostașii care erau în cetăți au fost uciși de către locuitori.
Pentru a prilejui celor fără gând rău o judecată isteață, omului tânăr cunoștință și bună cugetare. |
| Lyrics T. |
Pisti giuguri trecu anarga, lãi niori
Nicuchirã cu curunã: |
Peste juguri trec încet, nori negri
Ar trimite-o el acasă, |
| Small Corpus |
Unã caprã mushcrã shi unã ghesucãnutã pãshteau di unã parte, shi chipurle lã asunau: cing-cing.
Tsachilji giuca deavãrliga, arsãrea - tuts aleptsã, musheats, cu lãna lor ca mãtasea, di-lji yinea shi-al Tegã se-arsarã shi s-lu stringã-n bratsã picurarlu shi s-lu bashe, ahtare harauã lu-avea loatã! |
O capră cu botul bălțat și alta roșcată - pășteau răzlețite și tot le sunau tălăngile de la gât: cing-cing-cling.
Țapii se zbenguiau prin jur, țupăiau - frumoși toți, dragii, cu părul lor mătăsos, de-i venea lui Tegă, de bucurie, să-l ia în brațe pe cioban, să-l strângă la piept și să-l sărute, nestăpânit! |
| Phrases |
nu shtiu si mpartu palji la doi yumari
misuru stealili |
fi prost
hoinări |
| Avdhela |
Di toamnã-ascuturatã Ma,
s-ti caftu nji-easti fricã, |
Se scutură de toamnă,
Să te cer îmi este teamă, |
| Radio |
Deputatlu USR Cristian Seidler spusi cã proiectul di nom ari un impact multu modestu ti hãrgili bugetare cu pensiile spetsiali.
Timishoara fu singurã dinintea a represiuniljei trǎ ma multu di 5 dzãle di foc shi sãndze. |
Deputatul USR Cristian Seidler a afirmat că proiectul de lege are un impact foarte modest asupra cheltuielilor bugetare cu pensiile speciale.
Timișoara a fost singură în fața represiunii pentru cel puțin 5 zile de foc și sânge. |
| Aromanian Poetry |
Am un alt livend, di-Avdelǎ, alǎvdatlu-atsel di NUShI, te da njeate cãt sh-la-aushi;
Nã shcãmbã fãrã suflit i semn di tine, tutǎ shi salta siminatã di tine, tu ubor… |
În Avdela am un june tULLIU NUȘI - vorbă cerească versu-n veci o să-i trăiască !
O piatră doar aminte de tine mai aduce, și-o salcie sădită de mâna ta, cândva… |
| Tom Sawyer |
Arucã funea!
Mutri anvãrliga, canã… tu niheam di oarã u avea cartea tru mãnã. |
Dă și parâma!
Aruncă o privire de jur-împrejur, nu era nimeni… în clipa următoare, ținea cartea în mâini. |
| Divine Comedy |
Shi-agãrsheashti-n frixea tsi-lji da dãgoarea
tu doilji sherchi-mpriunats trupeashti |
pârjolul să-și ia ceva, căci la copil ia seamă
în cei doi șerpi împreunați |
| Modern poetry |
mashi suflitlu nu ari moarti
aclo iu suntu farurli apãryisiti. |
doar sufletul este fără moarte
acolo unde sunt farurile abandonate. |
| Tao Te Ching |
Pots s-ti tradz ditu minduirea a ta
Imnji dupã ea sh-nu-ari bitisitã. |
Poți să te dai înapoi din propria ta minte
urmează-l și nu are sfârșit. |
| The Little Prince |
Gioaca cor cathi gioi cu featili di-n hoara.
Elj avdu daima mashi alavdarli. |
Se duc să joace, joia, cu fetele din sat.
Vanitoșii nu aud niciodată decât laudele. |
| Writings Collection |
Li-alidzea ayonjia, unã dupu altã, ashi cum beai unã scafã cu apã!
Bãnarã deadunu tu Vãryãrie. |
Le citea repede, una după alta, așa cum ai bea un pahar cu apă!
Trăiseră împreună în Bulgaria. |
Our final data release contains an additional k dictionary pairs that we do not use to train the models. When training our initial machine translation models, we did not observe any improvements when using word-aligned dictionary entries. The entries are extracted from Papahagi’s Aromanian dictionary Papahagi (1963).
Appendix B Corpus statistics per source
| language | source | words | unique words | words/sentence |
| Aromanian | Bible | 493018 | 69780 | 16.15 |
| Lyrics T. | 16039 | 7586 | 3.98 | |
| Small corpus | 508090 | 74419 | 27.1 | |
| Phrases | 40941 | 9972 | 4.78 | |
| Avdhela | 7839 | 3567 | 3.48 | |
| Radio | 41036 | 12708 | 12.15 | |
| Aromanian Poetry | 36649 | 10433 | 11.69 | |
| Tom Sawyer | 23177 | 10325 | 13.13 | |
| Divine Comedy | 1269 | 747 | 4.83 | |
| Modern poetry | 11897 | 6336 | 5.3 | |
| Tao Te Ching | 5653 | 1886 | 2.62 | |
| The little prince | 3619 | 1625 | 5.81 | |
| Writings Collection | 17893 | 6646 | 12.81 | |
| Romanian | Bible | 580288 | 42487 | 19.0 |
| Lyrics T. | 16923 | 7022 | 4.2 | |
| Small corpus | 494098 | 53549 | 26.36 | |
| Phrases | 43590 | 7484 | 5.08 | |
| Avdhela | 8186 | 3330 | 3.63 | |
| Radio | 47588 | 12291 | 14.09 | |
| Aromanian Poetry | 38719 | 11126 | 12.35 | |
| Tom Sawyer | 24665 | 9296 | 13.97 | |
| Divine Comedy | 1397 | 744 | 5.31 | |
| Modern poetry | 14146 | 6412 | 6.31 | |
| Tao Te Ching | 4877 | 1740 | 2.26 | |
| The little prince | 3977 | 1717 | 6.38 | |
| Writings Collection | 19328 | 6398 | 13.84 | |
| English | Bible | 622430 | 28266 | 20.38 |
| Lyrics T. | 20277 | 5067 | 5.03 | |
| Small corpus | 518118 | 34576 | 27.64 | |
| Phrases | 49025 | 4950 | 5.72 | |
| Avdhela | 9466 | 2671 | 4.2 | |
| Radio | 53286 | 8348 | 15.78 | |
| Aromanian Poetry | 41069 | 7692 | 13.1 | |
| Tom Sawyer | 29411 | 6491 | 16.67 | |
| Divine Comedy | 1457 | 675 | 5.54 | |
| Modern poetry | 16959 | 4892 | 7.56 | |
| Tao Te Ching | 5089 | 1597 | 2.36 | |
| The little prince | 4258 | 1380 | 6.83 | |
| Writings Collection | 20998 | 4795 | 15.03 |
Appendix C LaBSE training hypermateres
| Training steps | 150000 |
| Batch size | 8 |
| Margin | 0.3 |
| Optimizer | Adafactor Shazeer and Stern (2018) |
| Learning rate | |
| Clip threshold |
Appendix D NLLB training hypermateres
| Training steps | 100000 |
| Batch size | 8 |
| Optimizer | Adafactor Shazeer and Stern (2018) |
| Scheduler | constant |
| Learning rate | |
| Weight Decay | |
| Clip threshold | |
| Maximum Sequence Length | 512 |
Appendix E LLM training
E.1 Hyperparameters
| Effective batch size | 32 |
| Number of epochs | 4 |
| Learning rate | |
| LR scheduler | cosine |
| Warmup steps | 100 |
| Weight decay | |
| Adam | |
| Adam | |
| Adam | |
| Maximum Sequence Length | 512 |
E.2 Prompt format
We format the prompts of the LLaMA 3.1 8B Instruct and RoLLaMA 3 8B Instruct models with the standard LLaMA 3.1 Instruct prompt format Dubey et al. (2024). For the Qwen and TowerInstruct models, we use their specific instruction prompt template Yang et al. (2024); Alves et al. (2024).
| System |
<|start_header_id|>system<|end_header_id|>
You are a helpful assistant capable of accurately translating between Aromanian, Romanian, and English. <|eot_id|> |
| User |
<|start_header_id|>user<|end_header_id|>
Translate this sentence from Romanian to Aromanian Adu-ți aminte că moartea nu zăbovește și hotărârea morții nu ți s-a arătat.<|eot_id|> |
| Model |
<|start_header_id|>assistant<|end_header_id|>
Adu-ts aminti cã moartea nu-amãnã shi-apofasea-a moartiljei nu tsã si-ari spusã. |
| System |
<|im_start|>system
You are a helpful assistant capable of accurately translating between Aromanian, Romanian, and English.<|im_end|> |
| User |
<|im_start|>user
Translate this sentence from Romanian to Aromanian Adu-ți aminte că moartea nu zăbovește și hotărârea morții nu ți s-a arătat.<|im_end|> |
| Model |
<|im_start|>assistant
Adu-ts aminti cã moartea nu-amãnã shi-apofasea-a moartiljei nu tsã si-ari spusã.<|im_end|> |
| User |
<|im_start|>user
Translate this sentence from Romanian to Aromanian Adu-ți aminte că moartea nu zăbovește și hotărârea morții nu ți s-a arătat.<|im_end|> |
| Model |
<|im_start|>assistant
Adu-ts aminti cã moartea nu-amãnã shi-apofasea-a moartiljei nu tsã si-ari spusã.<|im_end|> |
Appendix F Automatic evaluation results
| model | ron-¿rup | rup-¿ron | eng-¿rup | rup-¿eng | eng-¿ron | ron-¿eng |
| GPT-4o | 3.5 | 19.4 | 2.9 | 19.0 | 46.5 | 66.7 |
| Qwen2 7B Instruct | 15.2 | 28.8 | 10.3 | 32.4 | 35.4 | 62.9 |
| RoLLaMA | 14.7 | 29.7 | 11.0 | 32.7 | 38.1 | 64.4 |
| TowerInstruct v0.2 | 16.2 | 29.4 | 7.3 | 23.6 | 27.1 | 62.0 |
| LLaMA 3.1 8B Instruct | 15.8 | 30.2 | 10.7 | 33.7 | 38.6 | 65.1 |
| NLLB 1.3B | 16.2 | 31.9 | 12.8 | 35.6 | 41.4 | 63.0 |
| NLLB 600M | 17.0 | 30.9 | 13.4 | 34.9 | 39.1 | 62.2 |
| model | ron-¿rup | rup-¿ron | eng-¿rup | rup-¿eng | eng-¿ron | ron-¿eng |
| GPT-4o | 0.866 | 0.816 | 0.857 | 0.829 | 0.968 | 0.980 |
| Qwen2 7B Instruct | 0.905 | 0.896 | 0.880 | 0.903 | 0.955 | 0.976 |
| RoLLaMA 7B Instruct | 0.908 | 0.898 | 0.887 | 0.904 | 0.957 | 0.978 |
| TowerInstruct v0.2 | 0.904 | 0.894 | 0.840 | 0.870 | 0.923 | 0.973 |
| LLaMA 3.1 8B Instruct | 0.908 | 0.899 | 0.888 | 0.907 | 0.959 | 0.978 |
| NLLB 1.3B | 0.913 | 0.905 | 0.904 | 0.913 | 0.964 | 0.977 |
| NLLB 600M | 0.914 | 0.905 | 0.904 | 0.913 | 0.964 | 0.976 |